# Inclusive Training Separation and Implicit Knowledge Interaction for Balanced Online Class-Incremental Learning

### v1

@@ Abstract
Online class-incremental learning (OCIL) focuses on gradually learning new classes (called plasticity) from a stream of data in a single-pass, while concurrently preserving knowledge of previously learned classes (called stability). The primary challenge in OCIL lies in maintaining a good balance between the knowledge of old and new classes within the continually updated model. Most existing methods rely on explicit knowledge interaction through experience replay, and often employ exclusive training separation to address bias problems. Nevertheless, it still remains a big challenge to achieve a well-balanced learner, as these methods often exhibit either reduced plasticity or limited stability due to difficulties in continually integrating knowledge in the OCIL setting. In this paper, we propose a novel replay-based method, called Balanced Online Incremental Learning (BOIL), which can achieve both high plasticity and stability, thus ensuring more balanced performance in OCIL. Our BOIL method proposes an inclusive training separation strategy using dual classifiers so that knowledge from both old and new classes can effectively be integrated into the model, while introducing implicit approaches for transferring knowledge across the two classifiers. Extensive experimental evaluations over three widely-used OCIL benchmark datasets demonstrate the superiority of BOIL, showing more balanced yet better performance compared to state-of-the-art replay-based OCIL methods.
1 Introduction
Figure 1: Balance comparison in various OCIL methods on Split Mini-ImageNet ($\mathcal{M}$ = 1K). The color of the circle indicates the average accuracy. Lower forgetting and intransigence indicates better stability and higher plasticity, respectively. Our method (BOIL) is closest to the bottom-left corner with the highest accuracy, implying the most balanced and outperforming performance.
@@ 1 Introduction
Online class-incremental learning (OCIL) has attracted considerable attention in the deep learning community, enabling knowledge accumulation of new classes over time. Unlike its offline counterpart, which assumes access to a complete training set for each task, OCIL is given a single-pass stream where the learner is allowed to view each mini-batch of the task only once. This partial accessibility of data makes OCIL more appealing in practice but also brings a greater challenge in addressing the core problem of continual learning: stability-plasticity dilemma [1], i.e., balancing knowledge preservation (stability) with knowledge acquisition (plasticity). Due to the limited training opportunities with on-the-fly samples in OCIL, the model can easily be biased toward new classes without a careful strategy of knowledge preservation. Conversely, a strong policy to preserve past information can severely impair the learning performance due to the small number of incoming samples.
@@
The majority of OCIL methods tackle this challenge by using experience replay (ER) [2], which trains a mix batch of incoming samples from the stream and previous ones that are partially stored in a memory buffer. However, this explicit knowledge interaction via sample blending cannot completely resolve the stability-plasticity dilemma, since class imbalance is inevitable in OCIL with a fixed buffer size. As learning progresses over time, it would become more difficult to maintain balanced performance only by joint training with mix batches.
@@
To alleviate this imbalance issue in OCIL, existing replay-based methods incorporate their additional techniques, generally falling into two categories. The first category is training distinction [3, 4, 5, 6, 7, 8, 9], which imposes different training policies or losses on buffer samples and new samples, respectively. Most works in this category focus on how to avoid the trained model being biased toward new classes by loss separation [3, 4, 9]. However, this conversely tends to reduce the learning performance, as the conflicting tasks of learning new classes and replaying buffer samples are competitively and exclusively performed within a single classifier, without complementing each other. The second category is feature enhancement [10, 11, 12, 13, 14] that aims to obtain a more discriminative feature space where all embeddings, corresponding to either old classes or new classes, are well separated for better classification. This is typically done by using contrastive loss [11], as a replacement of cross-entropy loss, and by employing a post-hoc classifier, such as the nearest-class-mean (NCM) classifier [15], which takes the best use of well-separated features. However, in the OCIL setting with a limited number of samples, it is challenging to maintain historical feature information, as there are fewer pairs available for effective contrastive learning.
@@
To overcome these limitations, this paper proposes a novel replay-based OCIL method, called Balanced Online Incremental Learning (BOIL), designed to achieve a balanced yet satisfactory performance. Rather than competitively applying training policies or loss functions, our strategy is to incorporate separated components within the model itself, allowing for inclusive separation yet clearer differentiation across the tasks of knowledge acquisition and preservation. Specifically, we employ dual classifiers with identical structures: the stream classifier, which is dedicated to learning from new samples, and the buffer classifier, which focuses on preserving existing knowledge using buffered samples. This structural separation not only enables each classifier to concentrate on its corresponding task, but also facilitates learning knowledge across old and new classes, thereby enhancing knowledge acquisition and consolidation.
@@
For transferring knowledge across tasks, we introduce implicit techniques that enable effective knowledge exchange between the dual classifiers, beyond simply blending old and new samples. Our first approach is redesigning proxy-anchor loss (PAL) [16], so that the weights of the stream classifier are treated as if they are learnable proxies during training with the shared feature extractor. This method supports forward transfer by leveraging the previous knowledge of the feature extractor to guide the training of the stream classifier. For backward transfer, we employ a momentum-based updating mechanism that gradually transfers adaptive information from the stream classifier to the buffer classifier. With these techniques, both classifiers implicitly exchange and integrate their acquired knowledge.
@@
As summarized in Figure 1, our empirical study shows that the proposed BOIL method not only outperforms state-of-the-art OCIL methods in terms of overall performance (measured by average accuracy), but also clearly achieves the best balance between plasticity (measured by average intransigence) and stability (measured by average forgetting).
2 Related Works
Both online and offline continual learning (CL) typically focus on three major learning scenarios: class-incremental learning [17, 18, 19], task-incremental learning [18], and domain-incremental learning [20]. This paper specifically deals with online class-incremental learning (OCIL), which is practically appealing for real-world applications but more challenging due to the streaming nature of data.
Experience Replay.  Similar to the widely-used rehearsal method [21, 10, 22, 18] in offline CL, experience replay (ER) [2] is considered the most effective approach in OCIL. The baseline ER method utilizes a bounded memory buffer to store and replay historical samples while jointly learning from new samples. As ER inevitably encounters imbalance issues between the streaming and buffered samples due to the bounded buffer, existing OCIL methods have introduced additional techniques, such as training distinction and feature enhancement.
Training Distinction.  A number of existing works [6, 3, 23, 4, 9] can be classified as training distinction, which intends to apply different training policies or loss functions to buffer samples and stream samples to re-balance their contributions. For instance, [3] introduces loss separation, where losses for old and new classes are computed and backpropagated exclusively using a separated softmax function. Similarly, [4] proposes an asymmetric cross-entropy loss to avoid the drastic drift in old features caused by new incoming batches. While these approaches can help address imbalance issues mostly by prioritizing previous samples, their exclusive separation scheme within a single classifier can hinder the model’s ability to learn new knowledge as well as knowledge between old and new classes. Recently, [9] attempts to smooth this exclusive separation by carefully decoupling the cross-entropy loss for stream samples into two terms, one for only new classes and the other for old and new classes, while all the loss terms are still applied to a single classifier.
Feature Enhancement.  Another group [11, 13, 12, 14] focuses on feature enhancement, aiming to construct a more discriminative feature space by correcting or discarding biased features learned through cross-entropy loss. [11] leverages supervised contrastive loss for representation learning and employs the NCM classifier at inference time. [12] explores the coupling of proxy-based and contrastive-based loss by replacing anchor samples with proxies in contrastive loss. [13] introduces mutual information maximization with InfoNCE loss to overcome catastrophic forgetting in online CL. Similarly, [14] utilizes InfoNCE but introduces online prototype learning to obtain representative features. However, compared to cross-entropy loss, contrastive learning typically requires extensive sample comparisons, which is challenging in OCIL with the bounded buffer.
3 Problem Statement
Problem Setting.  In OCIL, we are given a sequence of tasks $\mathcal{D}=\{\mathcal{D}_{t}\}_{t=1}^{T}$ from a single-pass data stream. Each task $t$ is associated with its dataset $\mathcal{D}_{t}=\{\mathcal{X}_{t}\times\mathcal{Y}_{t}\}$, where $\mathcal{X}_{t}$ and $\mathcal{Y}_{t}$ represent samples and corresponding labels, respectively. Without overlapping classes across tasks, each task $t$ corresponds to a set $C_{t}$ of classes such that $|C_{t}|$ is the same for all tasks. For the replay method, we also maintain a bounded memory buffer $\mathcal{M}$ that stores some of the previously trained samples, which is therefore updated over the learning steps. At each learning step, the learner can access only a mini-batch $\mathcal{B}_{\mathcal{D}}\cup\mathcal{B}_{\mathcal{M}}$, where $\mathcal{B}_{\mathcal{D}}$ is a stream batch incoming from $\mathcal{D}$ and $\mathcal{B}_{\mathcal{M}}$ is a buffer batch retrieved from $\mathcal{M}$. Throughout the entire process, every instance in $\mathcal{D}$ can belong to only one stream batch $\mathcal{B}_{\mathcal{D}}$, and each $\mathcal{B}_{\mathcal{D}}$ can be trained only once.
@@ 3 Problem Statement
Figure 2: Training overview of our BOIL method. For each batch of stream and buffer samples, the shared feature extractor outputs corresponding features, which are then respectively fed into either stream or buffer classifier for separated training. Our redesigned proxy-anchor loss and momentum updating are used for implicit knowledge exchange, helping both forward and backward knowledge transfer and better discrimination ability. Green arrows represent the flow of features. Orange arrows indicate interactions between buffer features and the proxies from the stream classifier.
@@
$\underset{\Phi,\mathbf{W}}{\arg\min}\enspace\mathbb{E}_{(\mathbf{x},y)\sim\mathcal{D}_{t}\cup\mathcal{M}}\left[\mathcal{L}_{CE}\left(y,f(h(\mathbf{x};\Phi);\mathbf{W})\right)\right].$
@@
4 Methodology
In this section, we present our proposed BOIL method, which consists of two major components: (i) training separation with dual classifiers and (ii) implicit knowledge interaction. Intuitively, our techniques are motivated by both training distinction and feature enhancement, while introducing our novel approaches to address their limitations. The overall framework is illustrated in Figure 2.
4.1 Training Separation with Dual Classifiers
Figure 3: Comparison between two training separation schemes. (a) Exclusive separation [3] within a single classifier. (b) Inclusive separation with dual classifiers.
@@ 4.1 Training Separation with Dual Classifiers
Dual Classifiers for Training.  Based on the above motivation, our method proposes dual classifiers (DC) with the same structure, replacing the single classifier. These two independent classifiers, referred to as the stream classifier and the buffer classifier, are dedicated to the conflicting tasks of knowledge acquisition and preservation, respectively. Specifically, the stream classifier, parameterized by $\mathbf{W}_{str}$, primarily processes incoming batches (i.e., $\mathcal{B}_{\mathcal{D}}$) from the data stream, whereas the buffer classifier, parameterized by $\mathbf{W}_{buf}$, is fully responsible for learning from buffer batches (i.e., $\mathcal{B}_{\mathcal{M}}$). For the structure of each classifier, we adopt the idea of using cosine normalization [24, 25] that can eliminate the bias caused by difference in magnitudes, thereby alleviating imbalance. Formally, each classifier is defined as:
$f(\mathbf{z};\mathbf{W})=\eta\cdot\text{cos\_vec}(\mathbf{z},\mathbf{W}),$ (1)
[…]
@@
$\mathcal{L}_{DC}=\mathcal{L}_{CE_{str}^{\mathcal{D}}}+\alpha\mathcal{L}_{CE_{str}^{\mathcal{M}}}+(1-\alpha)\mathcal{L}_{CE_{buf}^{\mathcal{M}}},$ (2)
where the subscript $str$ or $buf$ refers to the stream or buffer classifier, and the superscript $\mathcal{D}$ or $\mathcal{M}$ indicates whether the loss is calculated based on the stream features $\mathbf{z}_{\mathcal{D}}$ or buffer features $\mathbf{z}_{\mathcal{M}}$. Note that this is a more refined version that can finely control the degree of separation by adjusting a control coefficient $\alpha$. When $\alpha=0$, it implies complete separation, whereas $\alpha\in(0,1)$ makes this separation smoother in a way that some buffer features can be fed into the stream classifier as well as the buffer classifier, which is indicated by the green dotted arrow in Figure 2. However, even in this smoother version, the buffer classifier is not designed to handle stream features for securing high stability.
@@
NCM Classifier for Inference.  After training the feature extractor with dual classifiers using $\mathcal{L}_{DC}$, we employ the NCM classifier for making inference. This is because NCM is known to be more robust against potential biases in trained classifiers, as noted by [11], as long as the feature extractor can generate well-discriminated feature embeddings. To be shown by our experimental results in Table 2, our BOIL method takes the best out of NCM possibly due to its well-trained feature extractor.
4.2 Implicit Knowledge Interaction

### v2

# Balanced Online Class-Incremental Learning via Dual Classifiers

@@ Abstract.
Online class-incremental learning (OCIL) focuses on gradually learning new classes (called plasticity) from a stream of data in a single-pass, while concurrently preserving knowledge of previously learned classes (called stability). The primary challenge in OCIL lies in maintaining a good balance between the knowledge of old and new classes within the continually updated model. Most existing methods rely on explicit knowledge interaction through experience replay, and often employ exclusive training separation to address bias problems. Nevertheless, it still remains a big challenge to achieve a well-balanced learner, as these methods often exhibit either reduced plasticity or limited stability due to difficulties in continually integrating knowledge in the OCIL setting. In this paper, we propose a novel replay-based method, called Balanced Inclusive Separation for Online iNcremental learning (BISON), which can achieve both high plasticity and stability, thus ensuring more balanced performance in OCIL. Our BISON method proposes an inclusive training separation strategy using dual classifiers so that knowledge from both old and new classes can effectively be integrated into the model, while introducing implicit approaches for transferring knowledge across the two classifiers. Extensive experimental evaluations over three widely-used OCIL benchmark datasets demonstrate the superiority of BISON, showing more balanced yet better performance compared to state-of-the-art replay-based OCIL methods.
Online Class-incremental Learning, Experience Replay, Continual Learning
journalyear: 2026copyright: ccconference: The 41st ACM/SIGAPP Symposium on Applied Computing; March 23–27, 2026; Thessaloniki, Greecebooktitle: The 41st ACM/SIGAPP Symposium on Applied Computing (SAC ’26), March 23–27, 2026, Thessaloniki, Greecedoi: 10.1145/3748522.3779821isbn: 979-8-4007-2294-3/2026/03ccs: Computing methodologies Machine learning
1. Introduction
@@
Online class-incremental learning (OCIL) has attracted considerable attention in the deep learning community, enabling knowledge accumulation of new classes over time. Unlike its offline counterpart, which assumes access to a complete training set for each task, OCIL is given a single-pass stream where the learner is allowed to view each mini-batch of the task only once. This partial accessibility of data makes OCIL more appealing in practice but also brings a greater challenge in addressing the core problem of continual learning: stability-plasticity dilemma (dilemma/mermillod2013stability), i.e., balancing knowledge preservation (stability) with knowledge acquisition (plasticity). Due to the limited training opportunities with on-the-fly samples in OCIL, the model can easily be biased toward new classes without a careful strategy of knowledge preservation. Conversely, a strong policy to preserve past information can severely impair the learning performance due to the small number of incoming samples.
Figure 1. Balance comparison in various OCIL methods on Split Mini-ImageNet ($\mathcal{M}$ = 1K). The color of the circle indicates the average accuracy. Lower forgetting and intransigence indicates better stability and higher plasticity, respectively. Our method (BISON) is closest to the bottom-left corner with the highest accuracy, implying the most balanced and outperforming performance.
@@
The majority of OCIL methods tackle this challenge by using experience replay (ER) (ER/chaudhry2019tiny), which trains a mix batch of incoming samples from the stream and previous ones that are partially stored in a memory buffer. However, this explicit knowledge interaction via sample blending cannot completely resolve the stability-plasticity dilemma, since class imbalance is inevitable in OCIL with a fixed buffer size. As learning progresses over time, it would become more difficult to maintain balanced performance only by joint training with mix batches.
@@
To alleviate this imbalance issue in OCIL, existing replay-based methods incorporate their additional techniques, generally falling into two categories. The first category is training distinction (SSIL/AhnKLBKM21; ACE/CacciaAATPB22; AFS/LiangCCJZ24; MIR/AljundiBTCCLP19; ASER/ShimMJSKJ21; GSS/AljundiLGB19; LODE/LiangL23), which imposes different training policies or losses on buffer samples and new samples, respectively. Most works in this category focus on how to avoid the trained model being biased toward new classes by loss separation (SSIL/AhnKLBKM21; ACE/CacciaAATPB22; LODE/LiangL23). However, this conversely tends to reduce the learning performance, as the conflicting tasks of learning new classes and replaying buffer samples are competitively and exclusively performed within a single classifier, without complementing each other. The second category is feature enhancement (iCaRL/RebuffiKSL17; SCR/MaiLKS21; PCR/LinZFLY23; OCM/GuoLZ22; OnPro/0001YHZS23) that aims to obtain a more discriminative feature space where all embeddings, corresponding to either old classes or new classes, are well separated for better classification. This is typically done by using contrastive loss (SCR/MaiLKS21), as a replacement of cross-entropy loss, and by employing a post-hoc classifier, such as the nearest-class-mean (NCM) classifier (NCM/MensinkVPC13), which takes the best use of well-separated features. However, in the OCIL setting with a limited number of samples, it is challenging to maintain historical feature information, as there are fewer pairs available for effective contrastive learning.
@@
To overcome these limitations, this paper proposes a novel replay-based OCIL method, called Balanced Inclusive Separation for Online iNcremental learning (BISON), designed to achieve a balanced yet satisfactory performance. Rather than competitively applying training policies or loss functions, our strategy is to incorporate separated components within the model itself, allowing for inclusive separation yet clearer differentiation across the tasks of knowledge acquisition and preservation. Specifically, we employ dual classifiers with identical structures: the stream classifier, which is dedicated to learning from new samples, and the buffer classifier, which focuses on preserving existing knowledge using buffered samples. This structural separation not only enables each classifier to concentrate on its corresponding task, but also facilitates learning knowledge across old and new classes, thereby enhancing knowledge acquisition and consolidation.
@@
For transferring knowledge across tasks, we introduce implicit techniques that enable effective knowledge exchange between the dual classifiers, beyond simply blending old and new samples. Our first approach is redesigning proxy-anchor loss (PAL) (PAL/KimKCK20), so that the weights of the stream classifier are treated as if they are learnable proxies during training. This method supports forward transfer by leveraging the previous knowledge of the feature extractor to guide the training of the stream classifier. For backward transfer, we employ a prototype-level proxy alignment feedback module that gradually transfers adaptive information from the stream classifier to the buffer classifier. With these techniques, both classifiers implicitly exchange and integrate their acquired knowledge.
@@
As summarized in Figure 1, our empirical study shows that the proposed BISON method not only outperforms state-of-the-art replay-based OCIL methods in terms of overall performance (average accuracy), but also clearly achieves the best balance between plasticity (average intransigence) and stability (average forgetting).
@@
Figure 2. Overview of our BISON method. For each batch of stream and buffer samples, the shared backbone outputs corresponding features $\mathbf{z}_{\mathcal{D}},\mathbf{z}_{\mathcal{M}}$, which are then respectively fed into either stream or buffer classifier $\mathbf{W}_{str/buf}$ for separated training. Our redesigned proxy-anchor loss $\mathcal{L}_{PAL}$ and proxy alignment feedback $\mathcal{L}_{Align}$ are used for implicit knowledge exchange, helping both forward and backward knowledge transfer and better discrimination ability. Arrows of different colors represent flow directions participating in different loss modules. Based on the labels of the buffer batch in the previous step, we only align the proxies of the corresponding classes in the buffer classifier to the stream classifier using stop-gradient (denoted by $\mathtt{sg}$). We use the whole buffer $\mathcal{M}$ to acquire the center of each class and the NCM classifier is applied for inference.
2. Related Works
Both online and offline continual learning (CL) typically focus on three major learning scenarios: class-incremental learning (WA/ZhaoXGZX20; DERPP/BuzzegaBPAC20; DER/YanX021), task-incremental learning (DERPP/BuzzegaBPAC20), and domain-incremental learning (three/VenTT22). This paper specifically deals with online class-incremental learning (OCIL), which is practically appealing for real-world applications but more challenging due to the streaming nature of data.
Experience Replay. Similar to the widely-used rehearsal method (GEM/Lopez-PazR17; iCaRL/RebuffiKSL17; MER/RiemerCALRTT19; DERPP/BuzzegaBPAC20) in offline CL, experience replay (ER) (ER/chaudhry2019tiny) is considered the most effective approach in OCIL. The baseline ER method utilizes a bounded memory buffer to store and replay historical samples while jointly learning from new samples. As ER inevitably encounters imbalance issues between the streaming and buffered samples due to the bounded buffer, existing OCIL methods have introduced additional techniques, such as training distinction and feature enhancement.
Training Distinction. A number of existing works (MIR/AljundiBTCCLP19; SSIL/AhnKLBKM21; DVC/Gu0WD22; ACE/CacciaAATPB22; LODE/LiangL23) can be classified as training distinction, which intends to apply different training policies or loss functions to buffer samples and stream samples to re-balance their contributions. For instance, (SSIL/AhnKLBKM21) introduces loss separation, where losses for old and new classes are computed and backpropagated exclusively using a separated softmax function. Similarly, (ACE/CacciaAATPB22) proposes an asymmetric cross-entropy loss to avoid the drastic drift in old features caused by new incoming batches. While these approaches can help address imbalance issues mostly by prioritizing previous samples, their exclusive separation scheme within a single classifier can hinder the model’s ability to learn new knowledge as well as knowledge between old and new classes. Recently, (LODE/LiangL23) attempts to smooth this exclusive separation by carefully decoupling the cross-entropy loss for stream samples into two terms, one for only new classes and the other for old and new classes, while all the loss terms are still applied to a single classifier.
Feature Enhancement. Another group (SCR/MaiLKS21; OCM/GuoLZ22; PCR/LinZFLY23; OnPro/0001YHZS23) focuses on feature enhancement, aiming to construct a more discriminative feature space by correcting or discarding biased features learned through cross-entropy loss. (SCR/MaiLKS21) leverages supervised contrastive loss for representation learning and employs the NCM classifier at inference time. (PCR/LinZFLY23) explores the coupling of proxy-based and contrastive-based loss by replacing anchor samples with proxies in contrastive loss. (OCM/GuoLZ22) introduces mutual information maximization with InfoNCE loss to overcome catastrophic forgetting in online CL. Similarly, (OnPro/0001YHZS23) utilizes InfoNCE but introduces online prototype learning to obtain representative features. However, compared to cross-entropy loss, contrastive learning typically requires extensive sample comparisons, which is challenging in OCIL with the bounded buffer.
3. Problem Statement
Problem Setting. In OCIL, we are given a sequence of tasks $\mathcal{D}=\{\mathcal{D}_{t}\}_{t=1}^{T}$ from a single-pass data stream. Each task $t$ is associated with its dataset $\mathcal{D}_{t}=\{\mathcal{X}_{t}\times\mathcal{Y}_{t}\}$, where $\mathcal{X}_{t}$ and $\mathcal{Y}_{t}$ represent samples and corresponding labels, respectively. Without overlapping classes across tasks, each task $t$ corresponds to a set $C_{t}$ of classes such that $|C_{t}|$ is the same for all tasks. For the replay method, we also maintain a bounded memory buffer $\mathcal{M}$ that stores some of the previously trained samples, which is therefore updated over the learning steps. At each learning step $s$, the learner can access only a mini-batch $\mathcal{B}^{s}_{\mathcal{D}}\cup\mathcal{B}^{s}_{\mathcal{M}}$, where $\mathcal{B}^{s}_{\mathcal{D}}$ is a stream batch incoming from $\mathcal{D}$ and $\mathcal{B}^{s}_{\mathcal{M}}$ is a buffer batch retrieved from $\mathcal{M}$. Throughout the entire process, every instance in $\mathcal{D}$ can belong to only one stream batch $\mathcal{B}_{\mathcal{D}}$, and each $\mathcal{B}_{\mathcal{D}}$ can be trained only once.
@@ 3. Problem Statement
$\underset{\Phi,\mathbf{W}}{\arg\min}\kern 5.0pt\mathbb{E}_{(\mathbf{x},y)\sim\mathcal{D}_{t}\cup\mathcal{M}}\left[\mathcal{L}_{CE}\left(y,f(h(\mathbf{x};\Phi);\mathbf{W})\right)\right].$
@@
4. Methodology
In this section, we present our proposed BISON method, which consists of two major components: (i) training separation with dual classifiers and (ii) implicit knowledge interaction. Intuitively, our techniques are motivated by both training distinction and feature enhancement, while introducing our novel approaches to address their limitations. The overall framework is illustrated in Figure 2.
Figure 3. Comparison between two training separation schemes. (a) Exclusive separation (SSIL/AhnKLBKM21) within a single classifier. (b) Inclusive separation with dual classifiers.
4.1. Training Separation with Dual Classifiers
@@
Dual Classifiers for Training. Based on the above motivation, our method proposes dual classifiers (DC) with the same structure, replacing the single classifier. These two independent classifiers, referred to as the stream classifier and the buffer classifier, are dedicated to the conflicting tasks of knowledge acquisition and preservation, respectively. Specifically, the stream classifier, parameterized by $\mathbf{W}_{str}$, primarily processes incoming batches (i.e., $\mathcal{B}_{\mathcal{D}}$) from the data stream, whereas the buffer classifier, parameterized by $\mathbf{W}_{buf}$, is fully responsible for learning from buffer batches (i.e., $\mathcal{B}_{\mathcal{M}}$). For the structure of each classifier, we adopt the idea of using cosine normalization (Cosine/LuoZXWRY18; Cosine/HouPLWL19) that can eliminate the bias caused by difference in magnitudes, thereby alleviating imbalance. Formally, the logit for the $j$-th class is defined as:
(1) $f_{j}(\mathbf{z};\mathbf{W})=\eta\cdot\text{cos}(\mathbf{W}_{j},\mathbf{z})$
where $f_{j}(\mathbf{z};\mathbf{W})$ is the resulting logit $s^{(j)}$ for the $j$-th class, $\mathbf{z}$ is the input feature vector, and $\mathbf{W}_{j}$ is the weight vector corresponding to the $j$-th class (i.e., the $j$-th row of the weight matrix $\mathbf{W}\in\mathbb{R}^{C\times D}$). The term $\eta$ is a trainable scaling factor, and the operator $\text{cos}(\mathbf{a},\mathbf{b})$ denotes the cosine similarity between two vectors $\mathbf{a}$ and $\mathbf{b}$, which is computed as $\frac{\mathbf{a}^{T}\mathbf{b}}{\|\mathbf{a}\|\|\mathbf{b}\|}$. As shown in Figure 2, both classifiers still share the feature extractor, from which we obtain the feature representations $\mathbf{z}_{\mathcal{D}}$ of incoming batches and those $\mathbf{z}_{\mathcal{M}}$ of buffer batches. Then, as indicated by the green solid arrows in Figure 2, $\mathbf{z}_{\mathcal{D}}$ and $\mathbf{z}_{\mathcal{M}}$ are fed into their respective classifiers to compute the logits and cross-entropy losses.
@@
(2) $\mathcal{L}_{DC}=\mathcal{L}_{CE_{str}^{\mathcal{D}}}+\alpha\mathcal{L}_{CE_{str}^{\mathcal{M}}}+(1-\alpha)\mathcal{L}_{CE_{buf}^{\mathcal{M}}},$
where the subscript strstr or bufbuf refers to the stream or buffer classifier, and the superscript $\mathcal{D}$ or $\mathcal{M}$ indicates whether the loss is calculated based on the stream features $\mathbf{z}_{\mathcal{D}}$ or buffer features $\mathbf{z}_{\mathcal{M}}$. The adaptive separation smoother $\alpha$ is designed as a task-specific learnable parameter, initialized anew for each task and processed via a sigmoid. Note that the $\alpha\in[0,1]$ and when $\alpha$ is closer to 0, it indicates stronger separation. This separation smoothing allows some buffer features to be fed into the stream classifier as well as the buffer classifier, which is indicated by the green dotted arrow in Figure 2. However, even in this smoother version, the buffer classifier is not designed to handle stream features for securing high stability.
@@
NCM Classifier for Inference. After training the feature extractor with dual classifiers using $\mathcal{L}_{DC}$, we employ the NCM classifier for making inference. This is because NCM is known to be more robust against potential biases in trained classifiers, as noted by (SCR/MaiLKS21), as long as the feature extractor can generate well-discriminated feature embeddings. To be shown by our experimental results in Table 3, our BISON method takes the best out of NCM possibly due to its well-trained feature extractor.
4.2. Implicit Knowledge Interaction